### v1

@@ 1 Introduction
A concept closely related to tree flow-sparsifiers is that of congestion-approximators. An $\alpha$-congestion-approximator of $G$ is a collection of cuts $\{(S_{i},V\setminus S_{i})\}_{i=1}^{k}$ that approximately capture the optimal congestion needed to route multi-commodity flows in $G$. Formally, for any demand matrix $Q$, consider the minimum ratio (over all cuts $(S_{i},V\setminus S_{i})$) between the capacity of the cut and the total demand separated by it, i.e., $\frac{\operatorname{cap}(S_{i},V\setminus S_{i})}{\operatorname{dem}_{Q}(S_{i},V\setminus S_{i})}$. The collection $\{(S_{i},V\setminus S_{i})\}$ is an $\alpha$-congestion-approximator if this ratio $\alpha$-approximates the minimum congestion required to route $Q$ in $G$. Tree flow-sparsifiers arising from hierarchical decompositions yield congestion approximators of the same quality, consisting of only a near-linear number of cuts. In a breakthrough result, Sherman [sherman2013nearly] showed that congestion-approximators of polylogarithmic quality suffice to compute approximate maximum flow in near-linear time.
@@
Räcke [racke2002minimizing] proved the existence of tree flow-sparsifiers of quality $O(\log^{3}n)$. The construction is based on a hierarchical decomposition of $G$. This result was soon improved to a polynomial time algorithm for computing a tree flow-sparsifier of quality $O(\log^{2}n\log\log n)$ [harrelson2003polynomial]. This series of works has culminated in an almost linear time (that is, $O\!\left(m^{1+o(1)}\right)$ time) construction of a tree flow-sparsifier of quality $O(\log^{4}n)$ by Räcke, Shah, and Täubig [racke2014computing], which was later improved to a near-linear time (that is, $\tilde{O}(m)$ time [^1] The $\tilde{O}$ notation is used to hide $\log^{c}n$ factors.) construction due to Peng’s near-linear time approximate max flow algorithm [peng2016approximate]. This $O(\log^{4}n)$ bound is the best known for near-linear time constructions of a tree flow-sparsifier and of congestion-approximators of near-linear size.
@@
Even though flow-sparsifiers are robust, for many applications a cut-sparsifier suffices (see, e.g., [chekuri2004all, engelberg2007cut, andreev2009simultaneous, bansal2014min, van2024almost]). The best known polynomial time construction by Räcke and Shah [RackeS14] gives a tree cut-sparsifier of quality $O(\log^{1.5}n\log\log n)$. Grid graphs impose an $\Omega(\log n)$ lower bound on the quality of tree cut-sparsifiers and flow-sparsifiers [bartal1997line, maggs1997exploiting].
@@
A tree flow-sparsifier $T$ of quality $\alpha$ is also a tree cut-sparsifier of the same quality. Therefore, [racke2014computing, peng2016approximate] implies a near-linear time construction of a tree cut-sparsifier of quality $O(\log^{4}n)$, and this is the best near-linear time construction to date. We give an $\tilde{O}(m)$ time construction of tree cut-sparsifier of quality $O(\log^{2}n\log\log n)$, a tree flow-sparsifier of quality $O(\log^{3}n\log\log n)$, and a congestion-approximator of near-linear size of quality $O(\log^{3}n\log\log n)$. This improves for the first time upon the 2014 bound of [racke2014computing, peng2016approximate]. The following theorem is our main result.
@@ 1.1 Techniques
We now elaborate on the tools and techniques that allows us to obtain Theorem 1.1. First, we develop a way to work directly with cuts rather than flows.
@@
For the cut-sparsifier setting, we replace explicit routings by a notion of $G$ respecting demand matrices. We say that $G$ $\alpha$-respects a demand matrix if for every cut $(S,\bar{S})$ the total amount of that demand that must cross $(S,\bar{S})$ is at most $\frac{1}{\alpha}\cdot\operatorname{cap}_{G}(S,\bar{S})$. To prove that $T$ is a tree cut-sparsifier we need to show that $G$ $\frac{1}{\alpha}$-respects every demand matrix that is $1$-respected by $T$. To carry out this task we use the notion of demand states (Definition 3.13). A demand state assigns to each vertex a vector of commodity masses, and each update to the state corresponds to moving mass. Demand states allow us to formally define and maintain invariants, reason about how mass is redistributed across clusters, and rigorously charge these movements to the capacity of cuts. In contrast to flow-sparsifiers, these movements and corresponding charges are not necessarily achieved by explicit routing. While similar techniques were used in previous work, using this demand state formulation allows us to provide a clearer and more systematic presentation of the algorithms, making their structure explicit.
@@
Expander Decomposition. Leveraging the “balanced cut or expander” primitive from [ADK25] with adaptively chosen expansion parameters, we develop a unified approach that allows us to simplify and strengthen the [racke2014computing] framework to obtain a tree cut-sparsifier of quality $O(\log^{3}n)$. Afterwards, we use the same approach to implement a variant of the refinement phase of [harrelson2003polynomial, RackeS14] in near-linear time, as we describe below.
@@
Using this primitive from [ADK25], together with the fair-cuts computation [LNPSsoda23], allows us to abstract away the details of the underlying algorithmic machinery and significantly simplify the construction of the tree cut-sparsifier, when compared to e.g., [racke2014computing]. This connects tree cut-sparsifier constructions to modern sparse-cut tools and shows how they can be constructed using these tools.
@@
A Near-Linear Refinement Phase. A main caveat in the construction of [racke2014computing] is that the boundary edges of the clusters are not flow-linked. The notion of a flow-linked set, introduced by [chekuri2005multicommodity], means that it is possible to route an all-to-all multi-commodity flow between the nodes (or edges) of this set. The lack of this property in [racke2014computing] adds a logarithmic factor to the quality of the resulting tree flow-sparsifier. For more details, see Section 2.4. In the context of tree cut–sparsifiers, flow-linkedness corresponds to expansion. By applying the “balanced cut or expander” primitive of [ADK25] to a given cluster $S$, we can test whether the boundary of $S$ expands. If not, we split $S$ along a balanced cut and recur on both sides. This recursive process, which we call the refinement phase, allows us to partition a cluster $S$ into $\{S_{1},\ldots,S_{t}\}$ such that in each $S_{i}$, the boundary edges expand.
@@
Additionally, similar to [harrelson2003polynomial, RackeS14], we show that there exists a low-congestion flow from the inter-cluster edges of the refinement partition to the boundary of $S$. We do this by carefully using the guarantees supplied by [ADK25]. This requires overcoming several technical obstacles, as we need to meticulously control both the number of boundary edges and the number of vertices in the resulting sub-clusters. Importantly, we are able to perform this step in near-linear time, instead of polynomial time.
@@
Previous constructions [harrelson2003polynomial, RackeS14] relied on intricate recursive schemes that handled “bad child events”, a subcall that violated the recursion invariants, by restarting parts of the computation. In contrast, our approach is entirely modular: we alternately apply a merge phase (on which we elaborate in Section 2.5) and a refinement phase, each performing a single, independent partition of the current cluster (no backtracking is needed). This simple alternation yields a much cleaner and modular tree cut-sparsifier construction of quality $O(\log^{2}n\log\log n)$.
@@ 1.2 Further Background
Our work builds on the tree flow-sparsifier constructions of [racke2014computing] and [harrelson2003polynomial], combined with an expander decomposition result by [ADK25].
@@
A crucial component in [racke2014computing] was a variant of the cut-matching game, introduced by Khandekar, Rao, and Vazirani [khandekar2009graph]. The cut-matching game [khandekar2009graph] provides an $O(\log^{2}n)$-approximate sparsest cut in near-linear time. By running a “non-stop” version of this game, [racke2014computing] obtained balanced $O(\log^{2}n)$-approximate sparse cuts, or alternatively, constructed clusters whose inter-cluster edges (not including boundary edges) are flow-linked.
@@
A stronger cut strategy was developed by [orecchia2008partitioning], achieving an $O(\log n)$-approximation for the sparsest cut problem. The expansion produced by this cut strategy follows from Cheeger’s inequality [cheeger1970lower] and therefore, at least intuitively, is unsuitable for flow routing purposes. This stands in contrast to the cut player of [khandekar2009graph], whose expansion follows from embedding a uniform demands matrix (all $1$’s) into the expanding set.
@@
Agassy, Dorfman, and Kaplan [agassy2023expander] were able to translate [orecchia2008partitioning]’s improved cut strategy into a “balanced cut or expander” procedure of approximation $O(\log n)$, resulting in a near-linear time expander decomposition algorithm which improved on the prior expander decomposition algorithm by Saranurak and Wang [saranurak2019expander] that uses [khandekar2009graph]’s cut strategy. In [ADK25], Agassy, Dorfman and Kaplan show how to extend their results to the case of generalized conductance. By generalized conductance, they refer to conductance with respect to a vertex measure. Formally, given a vertex measure $\mu:V\to\mathop{\mathbb{R}}_{\geq 0}$, and a set $S\subseteq V$, the expansion of a cut $(S,\bar{S})$ with respect to $\mu$ is defined as $\Phi^{\mu}_{G}(S,\bar{S})\mathrel{\mathop{:}}=\frac{\operatorname{cap}(S,\bar{S})}{\min(\mu(S),\mu(\bar{S}))}$, where $\mu(U)$, for $U\subseteq V$, is defined as $\mu(U)\mathrel{\mathop{:}}=\sum_{v\in U}{\mu(v)}$. We use the result by [ADK25] in a black-box manner as explained in Section 2.5.
@@
Typically, as in this work, hierarchical decompositions are computed top-down by recursively splitting clusters starting from the root $V$. Li, Rao, and Wang [li2025congestion] take the opposite, bottom-up approach, constructing a congestion approximator of quality $O(\log^{10}n)$. Their main tool is a weak expander decomposition: they build a sequence of partitions $P_{1},P_{2},\ldots,P_{L}$ of $V$, where each partition corresponds to a weak expander decomposition, and $P_{1}$ is the partition into singleton clusters and $P_{L}=\{V\}$. The collection of cuts from these partitions does not form a laminar family. Instead, the congestion-approximator’s set of cuts consists of all intersections of sub-clusters drawn from different levels. A technical challenge is that expander-decomposition routines typically invoke approximate max-flow algorithms that themselves rely on the top-down congestion-approximator of [racke2014computing]. The construction of [li2025congestion] avoids this potential circularity by maintaining “pseudo”-congestion-approximators derived from the partial hierarchy $P_{1},\ldots,P_{i}$ and using them within the max-flow subroutines required for the weak expander decomposition.
@@
We note that the “balanced cut or expander” primitive of [ADK25], which we use extensively, internally computes fair-cuts (approximate maximum flows) via the algorithm of [LNPSsoda23]. This algorithm requires a congestion–approximator for the underlying graph $G$. By instantiating the congestion–approximator required by the fair-cut routine [LNPSsoda23] with the construction of [li2025congestion], our tree cut–sparsifiers can be implemented without invoking the congestion–approximator of [racke2014computing]. This means that one can also think of our work (in particular, Sections 4 and 5) as a simplification of the construction of [racke2014computing] (which does not depend on it).
@@ 1.3 Related Work
When allowing to sparsify with a convex combination of trees, Räcke achieved a polynomial time algorithm for computing a flow-sparsifier [racke2008optimal] of optimal $O(\log n)$ quality. However, some problems require a flow-sparsifier which is a single tree (e.g., [khandekar2014advantage, andreev2009simultaneous, bansal2014min]).
Using modern techniques of (dynamic) expander decomposition, and expander-hierarchies, [goranci2021expander] achieved a fully dynamic, $n^{o(1)}$ update and $O(\log^{1/6}n)$ query time, data structure for maintaining a tree flow-sparsifier of quality $n^{o(1)}$.
Tree flow–sparsifiers have been extensively used in the context of oblivious routing schemes, that is, routings that are independent of the input demands and preassign flow paths for every pair of vertices. There has also been progress on hop-constrained routing, where paths are kept short. Ghaffari, Haeupler, and Žužić obtain a static oblivious scheme that, for any hop bound $h$, uses paths of length at most $h\cdot\mathrm{poly}(\log n)$ and achieves congestion within $n^{o(1)}$ of the best routing that uses paths of length at most $h$ [ghaffari2021hop]. Haeupler, Räcke, and Ghaffari introduce hop-constrained expander decompositions, from which they derive hop-constrained routing [haeupler2022hop]. These works are complementary to our goal of producing a single tree with provable cut guarantees and near-linear preprocessing time.
@@ 2.1 Preliminaries
We use $\log$ to denote the base-2 logarithm. Throughout the paper, we assume all graphs are undirected. Given an undirected graph $G=(V,E)$ and a set $S\subseteq V$, we denote by $\deg_{G}(v)$ the degree of $v$ (for a weighted graph, this is the sum of the weights of the edges adjacent to $v$). The volume of a set $S$ is defined as $\operatorname{\mathbf{vol}}_{G}(S)\mathrel{\mathop{:}}=\sum_{v\in S}\deg_{G}(v)$. Given vertex weights $\mu:V\to\mathop{\mathbb{R}}_{\geq 0}$, and a set $U\subseteq V$, we denote the total weight of $U$ as $\mu(U)\mathrel{\mathop{:}}=\sum_{v\in U}{\mu(v)}$. In almost all of our use cases, $\mu$ will be an indicator measure. That is, $\mu(v)=\mathbbm{1}_{\{v\in X\}}$ for some $X\subseteq V$, and $\mu$ simply distinguishes the set $X$ of vertices (formally, $\mu(v)=1$ if $v\in X$ and otherwise $\mu(v)=0$). For disjoint sets $A,B\subseteq V$, we denote by $E_{G}(A,B)$ the set of edges connecting $A$ and $B$. We denote by $\operatorname{cap}_{G}(A,B)$ the number of edges in $E_{G}(A,B)$, or the sum of their weights if the graph is weighted. We sometimes omit the subscript when the graph is clear from the context. Denote $\bar{A}=V\setminus A$. If $A$ and $\bar{A}$ are both non-empty, then we call $(A,\bar{A})$ a cut.
@@
We identify a multi-commodity flow problem with its demand matrix, and use $Q$ and $D_{Q}$ interchangeably.
@@
Let $G=(V,E)$ and let $A\subseteq V$. The all-to-all flow problem on $A$ is defined by the demand matrix $D_{A}(u,v)\mathrel{\mathop{:}}=\begin{cases*}\frac{1}{|A|},&if $u,v\in A$,\\ 0,&otherwise.\end{cases*}$.
@@
Given a graph $G=(V,E)$ with edge capacities $c:E\to\mathbb{R}_{>0}$ and a flow $f:E\to\mathbb{R}_{\geq 0}$ routing a multi-commodity flow problem $Q$ ($f$ does not necessarily abide capacity constraints), the congestion of $f$ in $G$ is
@@
We say that $H$ is a cut-sparsifier of $G$ of quality $\alpha$ if for every $S\subseteq V$ it holds that
@@
We say that $H$ is a flow-sparsifier of $G$ of quality $\alpha$ if for every multi-commodity flow problem $Q$ on the vertex set $V$, it holds that
@@
Let $Q$ be a multi-commodity flow problem on $G=(V,E)$, with demand matrix $D_{Q}$. Let $(S,\bar{S})$ be a cut in $G$.
@@
Note that a $G$ $\alpha$-respecting a demand matrix $Q$ does not necessarily imply that $Q$ can be routed with congestion $1/\alpha$ in $G$. Due to the flow–cut gap [leighton1999multicommodity], an additional $O(\log n)$ factor may be required.
@@
If the multi-commodity flow problem $Q$ can be routed in $G$ with congestion $\alpha$, then $\operatorname{dem}_{Q}(B,W)\leq\alpha\cdot\operatorname{cap}_{G}(S,V\setminus S)$, for all cuts $(S,V\setminus S)$, i.e., $G$ $\frac{1}{\alpha}$-respects $Q$.

### v2

@@ 1 Introduction
A concept closely related to tree flow-sparsifiers is that of congestion-approximators. An $\alpha$-congestion-approximator of $G$ is a collection of cuts $\{(S_{i},V\setminus S_{i})\}_{i=1}^{k}$ that approximately capture the optimal congestion needed to route multi-commodity flows in $G$. Formally, for any demand matrix $Q$, consider the minimum ratio (over all cuts $(S_{i},V\setminus S_{i})$) between the capacity of the cut and the total demand separated by it, i.e., $\frac{\operatorname{cap}(S_{i},V\setminus S_{i})}{\operatorname{dem}_{Q}(S_{i},V\setminus S_{i})}$. The collection $\{(S_{i},V\setminus S_{i})\}$ is an $\alpha$-congestion-approximator if this ratio $\alpha$-approximates the minimum congestion required to route $Q$ in $G$. Tree flow-sparsifiers arising from hierarchical decompositions yield congestion approximators of the same quality, consisting of only a near-linear number of cuts. In a breakthrough result, Sherman [33] showed that congestion-approximators of polylogarithmic quality suffice to compute approximate maximum flow in near-linear time.
@@
Räcke [29] proved the existence of tree flow-sparsifiers of quality $O(\log^{3}n)$. The construction is based on a hierarchical decomposition of $G$. This result was soon improved to a polynomial time algorithm for computing a tree flow-sparsifier of quality $O(\log^{2}n\log\log n)$ [15]. This series of works has culminated in an almost linear time (that is, $O\!\left(m^{1+o(1)}\right)$ time) construction of a tree flow-sparsifier of quality $O(\log^{4}n)$ by Räcke, Shah, and Täubig [27], which was later improved to a near-linear time (that is, $\tilde{O}(m)$ time [^1] The $\tilde{O}$ notation is used to hide $\log^{c}n$ factors.) construction due to Peng’s near-linear time approximate max flow algorithm [26]. This $O(\log^{4}n)$ bound is the best known for near-linear time constructions of a tree flow-sparsifier and of congestion-approximators of near-linear size.
@@
Even though flow-sparsifiers are robust, for many applications a cut-sparsifier suffices (see, e.g., [9, 11, 3, 5, 34]). The best known polynomial time construction by Räcke and Shah [28] gives a tree cut-sparsifier of quality $O(\log^{1.5}n\log\log n)$. Grid graphs impose an $\Omega(\log n)$ lower bound on the quality of tree cut-sparsifiers and flow-sparsifiers [6, 23].
@@
A tree flow-sparsifier $T$ of quality $\alpha$ is also a tree cut-sparsifier of the same quality. Therefore, [27, 26] implies a near-linear time construction of a tree cut-sparsifier of quality $O(\log^{4}n)$, and this is the best near-linear time construction to date. We give an $\tilde{O}(m)$ time construction of tree cut-sparsifier of quality $O(\log^{2}n\log\log n)$, a tree flow-sparsifier of quality $O(\log^{3}n\log\log n)$, and a congestion-approximator of near-linear size of quality $O(\log^{3}n\log\log n)$. This improves for the first time upon the 2014 bound of [27, 26]. The following theorem is our main result.
@@ 1.1 Techniques
We now elaborate on the tools and techniques that allow us to obtain Theorem 1.1. First, we develop a way to work directly with cuts rather than flows.
@@
For the cut-sparsifier setting, we replace explicit routings by a notion of $G$ respecting demand matrices. We say that $G$ $\alpha$-respects a demand matrix if for every cut $(S,\bar{S})$ the total amount of that demand that must cross $(S,\bar{S})$ is at most $\frac{1}{\alpha}\cdot\operatorname{cap}_{G}(S,\bar{S})$. To prove that $T$ is a tree cut-sparsifier we need to show that $G$ $\frac{1}{\alpha}$-respects every demand matrix that is $1$-respected by $T$ (see Claim 5.2). To carry out this task we use the notion of demand states (Definition 3.12). A demand state assigns to each vertex a vector of commodity masses, and each update to the state corresponds to moving mass. Demand states allow us to formally define and maintain invariants, reason about how mass is redistributed across clusters, and rigorously charge these movements to the capacity of cuts. In contrast to flow-sparsifiers, these movements and corresponding charges are not necessarily achieved by explicit routing. While similar techniques were used in previous work, using this demand state formulation allows us to provide a clearer and more systematic presentation of the algorithms, making their structure explicit.
@@
Expander Decomposition. Leveraging the “balanced cut or expander” primitive from [2] with adaptively chosen expansion parameters, we develop a unified approach that allows us to simplify and strengthen the [27] framework to obtain a tree cut-sparsifier of quality $O(\log^{3}n)$. Afterwards, we use the same approach to implement a variant of the refinement phase of [15, 28] in near-linear time, as we describe below.
@@
Using this primitive from [2], together with the fair-cuts computation [21], allows us to abstract away the details of the underlying algorithmic machinery and significantly simplify the construction of the tree cut-sparsifier, when compared to e.g., [27]. This connects tree cut-sparsifier constructions to modern sparse-cut tools and shows how they can be constructed using these tools.
@@
A Near-Linear Time Refinement Phase. A main caveat in the construction of [27] is that the boundary edges of the clusters are not flow-linked. The notion of a flow-linked set, introduced by [10], means that it is possible to route an all-to-all multi-commodity flow between the nodes (or edges) of this set. The lack of this property in [27] adds a logarithmic factor to the quality of the resulting tree flow-sparsifier. For more details, see Section 2.4. In the context of tree cut–sparsifiers, flow-linkedness corresponds to expansion. By applying the “balanced cut or expander” primitive of [2] to a given cluster $S$, we can test whether the boundary of $S$ expands. If not, we split $S$ along a balanced cut and recur on both sides. This recursive process, which we call the refinement phase, allows us to partition a cluster $S$ into $\{S_{1},\ldots,S_{t}\}$ such that in each $S_{i}$, the boundary edges expand.
@@
Additionally, similar to [15, 28], we show that there exists a low-congestion flow from the inter-cluster edges of the refinement partition to the boundary of $S$. We do this by carefully using the guarantees supplied by [2]. This requires overcoming several technical obstacles, as we need to meticulously control both the number of boundary edges and the number of vertices in the resulting sub-clusters. Importantly, we are able to perform this step in near-linear time, instead of polynomial time.
@@
Previous constructions [15, 28] relied on intricate recursive schemes that handled “bad child events”, a subcall that violated the recursion invariants, by restarting parts of the computation. In contrast, our approach is entirely modular: we alternately apply a merge phase (on which we elaborate in Section 2.5) and a refinement phase, each performing a single, independent partition of the current cluster (no backtracking is needed). This simple alternation yields a much cleaner and modular tree cut-sparsifier construction of quality $O(\log^{2}n\log\log n)$.
@@ 1.2 Further Background
Our work builds on the tree flow-sparsifier constructions of [27] and [15], combined with an expander decomposition result by [2].
@@
A crucial component in [27] was a variant of the cut-matching game, introduced by Khandekar, Rao, and Vazirani [18]. The cut-matching game [18] provides an $O(\log^{2}n)$-approximate sparsest cut in near-linear time. By running a “non-stop” version of this game, [27] obtained balanced $O(\log^{2}n)$-approximate sparse cuts, or alternatively, constructed clusters whose inter-cluster edges (not including boundary edges) are flow-linked.
@@
A stronger cut strategy was developed by [25], achieving an $O(\log n)$-approximation for the sparsest cut problem. The expansion produced by this cut strategy follows from Cheeger’s inequality [8] and therefore, at least intuitively, is unsuitable for flow routing purposes. This stands in contrast to the cut player of [18], whose expansion follows from embedding a uniform demands matrix (all $1$’s) into the expanding set.
@@
Agassy, Dorfman, and Kaplan [1] were able to translate [25]’s improved cut strategy into a “balanced cut or expander” procedure of approximation $O(\log n)$, resulting in a near-linear time expander decomposition algorithm which improved on the prior expander decomposition algorithm by Saranurak and Wang [32] that uses [18]’s cut strategy. In [2], Agassy, Dorfman and Kaplan show how to extend their results to the case of generalized conductance. By generalized conductance, they refer to conductance with respect to a vertex measure. Formally, given a vertex measure $\mu:V\to\mathop{\mathbb{R}}_{\geq 0}$, and a set $S\subseteq V$, the expansion of a cut $(S,\bar{S})$ with respect to $\mu$ is defined as $\Phi^{\mu}_{G}(S,\bar{S})\mathrel{\mathop{:}}=\frac{\operatorname{cap}(S,\bar{S})}{\min(\mu(S),\mu(\bar{S}))}$, where $\mu(U)$, for $U\subseteq V$, is defined as $\mu(U)\mathrel{\mathop{:}}=\sum_{v\in U}{\mu(v)}$. We use the result by [2] in a black-box manner as explained in Section 2.5. We note that for the purpose of obtaining our $O(\log^{3}\log\log n)$ quality tree flow-sparsifier, one could alternatively rely on [32], together with our refinement phase. However, achieving the improved $O(\log^{2}\log\log n)$ quality tree cut-sparsifier crucially relies on the spectral cut player underlying [2].
@@
Typically, as in this work, hierarchical decompositions are computed top-down by recursively splitting clusters starting from the root $V$. Li, Rao, and Wang [22] take the opposite, bottom-up approach, constructing a congestion approximator of quality $O(\log^{10}n)$. Their main tool is a weak expander decomposition: they build a sequence of partitions $P_{1},P_{2},\ldots,P_{L}$ of $V$, where each partition corresponds to a weak expander decomposition, and $P_{1}$ is the partition into singleton clusters and $P_{L}=\{V\}$. The collection of cuts from these partitions does not form a laminar family. Instead, the congestion-approximator’s set of cuts consists of all intersections of sub-clusters drawn from different levels. A technical challenge is that expander-decomposition routines typically invoke approximate max-flow algorithms that themselves rely on the top-down congestion-approximator of [27]. The construction of [22] avoids this potential circularity by maintaining “pseudo”-congestion-approximators derived from the partial hierarchy $P_{1},\ldots,P_{i}$ and using them within the max-flow subroutines required for the weak expander decomposition.
@@
We note that the “balanced cut or expander” primitive of [2], which we use extensively, internally computes fair-cuts (approximate maximum flows) via the algorithm of [21]. This algorithm requires a congestion–approximator for the underlying graph $G$. By instantiating the congestion–approximator required by the fair-cut routine [21] with the construction of [22], our tree cut–sparsifiers can be implemented without invoking the congestion–approximator of [27]. This means that one can also think of our work (in particular, Sections 4 and 5) as a simplification of the construction of [27] (which does not depend on it).
@@ 1.3 Related Work
When allowing to sparsify with a convex combination of trees, Räcke achieved a polynomial time algorithm for computing a flow-sparsifier [30] of optimal $O(\log n)$ quality. However, some problems require a flow-sparsifier which is a single tree (e.g., [17, 3, 5]).
Using modern techniques of (dynamic) expander decomposition, and expander-hierarchies, [13] achieved a fully dynamic, $n^{o(1)}$ update and $O(\log^{1/6}n)$ query time, data structure for maintaining a tree flow-sparsifier of quality $n^{o(1)}$.
Tree flow–sparsifiers have been extensively used in the context of oblivious routing schemes, that is, routings that are independent of the input demands and preassign flow paths for every pair of vertices. There has also been progress on hop-constrained routing, where paths are kept short. Ghaffari, Haeupler, and Žužić obtain a static oblivious scheme that, for any hop bound $h$, uses paths of length at most $h\cdot\mathrm{poly}(\log n)$ and achieves congestion within $n^{o(1)}$ of the best routing that uses paths of length at most $h$ [12]. Haeupler, Räcke, and Ghaffari introduce hop-constrained expander decompositions, from which they derive hop-constrained routing [14]. These works are complementary to our goal of producing a single tree with provable cut guarantees and near-linear preprocessing time.
Recent Work.
In an independent and parallel work Henzinger, Münk, and Räcke [16] achieved a tree cut sparsifier with the same $O(\log^{2}n\log\log n)$ quality. They also show how to implement their construction in parallel with polylogarithmic span and near linear work. Their hierarchy construction follows the [15] bad-child-event paradigm: when a cluster cannot be certified to satisfy the required expansion invariant, the algorithm dynamically restructures the current hierarchy level and revisits the affected subproblems. This event-driven backtracking leads to a more intricate control flow. In contrast, our construction avoids such restart mechanisms entirely. The merge and refinement phases are applied as independent, one-shot procedures, resulting in a simpler and more modular algorithmic structure.
@@ 2.1 Preliminaries
We use $\log$ to denote the base-2 logarithm. Throughout the paper, we assume all graphs are undirected. Given an undirected graph $G=(V,E)$ and a set $S\subseteq V$, we denote by $\deg_{G}(v)$ the degree of $v$ (for a weighted graph, this is the sum of the weights of the edges adjacent to $v$). The volume of a set $S$ is defined as $\operatorname{\mathbf{vol}}_{G}(S)\mathrel{\mathop{:}}=\sum_{v\in S}\deg_{G}(v)$. Given vertex weights $\mu:V\to\mathop{\mathbb{R}}_{\geq 0}$, and a set $U\subseteq V$, we denote the total weight of $U$ as $\mu(U)\mathrel{\mathop{:}}=\sum_{v\in U}{\mu(v)}$. We refer to $\mu$ interchangeably as a vertex weighting or a vertex measure. In almost all of our use cases, $\mu$ will be an indicator measure. That is, $\mu(v)=\mathbbm{1}_{\{v\in X\}}$ for some $X\subseteq V$, and $\mu$ simply distinguishes the set $X$ of vertices (formally, $\mu(v)=1$ if $v\in X$ and otherwise $\mu(v)=0$). For disjoint sets $A,B\subseteq V$, we denote by $E_{G}(A,B)$ the set of edges connecting $A$ and $B$. We denote by $\operatorname{cap}_{G}(A,B)$ the number of edges in $E_{G}(A,B)$, or the sum of their weights if the graph is weighted. We sometimes omit the subscript when the graph is clear from the context. Denote $\bar{A}=V\setminus A$. If $A$ and $\bar{A}$ are both non-empty, then we call $(A,\bar{A})$ a cut.
@@
We identify a multi-commodity flow problem with its demand matrix (or simply demands), and use $Q$ and $D_{Q}$ interchangeably.
@@
Let $G=(V,E)$ and let $A\subseteq V$. The all-to-all multi-commodity flow problem on $A$ is defined by the demand matrix $D_{A}(u,v)\mathrel{\mathop{:}}=\begin{cases*}\frac{1}{|A|},&if $u,v\in A$,\\ 0,&otherwise.\end{cases*}$.
@@
Given a graph $G=(V,E)$ with edge capacities $c:E\to\mathbb{R}_{>0}$ and a flow $f:E\to\mathbb{R}_{\geq 0}$ routing a demand matrix $Q$ ($f$ does not necessarily abide capacity constraints), the congestion of $f$ in $G$ is
@@
We say that $H=(V_{H},E_{H})$ is a cut-sparsifier of $G=(V,E)$ of quality $\alpha$ if $V\subseteq V_{H}$ and for every $S\subseteq V$ it holds that
@@
Given a graph $G=(V,E)$, we say that a graph $H$ on the vertices $V_{H}$, where $V\subseteq V_{H}$, is a flow-sparsifier of $G$ of quality $\alpha$ if for every demand matrix $Q$ on the vertex set $V$, it holds that
@@
Let $Q$ be a demand matrix on $G=(V,E)$ and let $(S,\bar{S})$ be a cut in $G$.
@@
[…]
@@
If the demand matrix $Q$ can be routed in $G$ with congestion $\alpha$, then $\operatorname{dem}_{Q}(S,V\setminus S)\leq\alpha\cdot\operatorname{cap}_{G}(S,V\setminus S)$, for all cuts $(S,V\setminus S)$, i.e., $G$ $\frac{1}{\alpha}$-respects $Q$.
Note that $G$ $\alpha$-respecting a demand matrix $Q$ does not necessarily imply that $Q$ can be routed with congestion $1/\alpha$ in $G$. Due to the flow–cut gap [19], an additional $O(\log n)$ factor may be required.